OpenStudy (anonymous):

What is sin(pi/3 + x) - sin(pi/3 - x) equal to ?

OpenStudy (anonymous):

Help!!!!

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Trig_Identity#Angle_sum_and_difference_identities use the sum/difference identities

OpenStudy (dumbcow):

\[(\sin(\pi/3)\cos x + \cos(\pi/3) \sin x) - (\sin(\pi/3)\cos x - \cos(\pi/3) \sin x)\]

OpenStudy (anonymous):

Yes but then how do you simplify ?

OpenStudy (anonymous):

Sry thanks a lot, was stupid

OpenStudy (dumbcow):

haha ok you got it :) = sin x